OpenStudy (anonymous):

What is the solution to the equation? 4.8+x-5.9 = 3.9

OpenStudy (anonymous):

Well this is just simple algebra, to do this what you must do to one side you must also do to the other. For example, lets take this equation. x+4.2=4 So we want to get x on its own to do this, we want to get rid of 4.2 on the left side. What operation gets rid of 4.2? Lets try the opposite of adding, which is subtracting, so we subtract 4.2 on both sides and it gets cancelled on the left and we're left with x = 4-4.2, which gives us x = -0.2 Now try this with your equation and see what you get.